# The Sparse Frontier: Sparse Attention Trade-offs in Transformer LLMs

### v1

@@ 3.1 Models
We perform experiments on Qwen 2.5 models (Yang et al., 2024a) (7B, 14B, 32B, 72B parameters). This model family was selected to systematically investigate how model size interacts with sequence length, task characteristics, and sparsity patterns. Qwen 2.5 uniquely supports 128k context length and provides multiple model sizes trained with consistent methodology—critical for controlled scaling experiments. We modify only the attention mechanism, preserving the original architectures, and utilise the vLLM inference engine (Kwon et al., 2023) with full bf16 precision. Implementation details for each sparse attention pattern are in Section A.1.
@@ 3.2 Sparse Attention Methods
We evaluate a series of sparse attention methods, which we choose as a representative set spanning across the key dimensions described in Section 2. We focus exclusively on content-aware methods, as prior work has demonstrated that fixed patterns consistently underperform their content-aware counterparts (Li et al., 2025b): the full list is presented in Table 1.
@@ 3.3 Tasks
We evaluate 9 diverse tasks selected to reflect different characteristics along 3 key dimensions known to influence sparse attention performance: task difficulty—defined by Dispersion (how hard it is to locate necessary information) and Scope (how much information must be processed) (Goldman et al., 2024)—and data Naturalness (natural language vs. synthetic data). This multi-dimensional approach is motivated by recent findings that attention patterns vary significantly across task types: retrieval tasks often exhibit localised attention, while reasoning tasks show more uniform distributions that are challenging for sparse methods (Liu et al., 2025c; Chen et al., 2024; Li et al., 2025b). The naturalness dimension is also crucial, as synthetic tasks with arbitrary symbol sequences yield different token representation distributions compared to natural language (Liu et al., 2024). Our task suite therefore incorporates four core tasks from the RULER benchmark (Hsieh et al., 2024)—Retrieval (NIAH), Multi-hop reasoning (VT), Aggregation (CWE), and QA (SQuAD)—to provide controlled environments (mostly synthetic) for specific capabilities. We complement these with natural text tasks from existing benchmarks assessed for minimal contamination risk (Li et al., 2024a), such as QA from QuALITY and TOEFL; however these are low-dispersion, low-scope tasks. Thus, we additionally introduce three novel tasks (Story Retrieval, Multi-hop, Filtering) that translate RULER’s challenging tasks (with high dispersion or scope) into naturalistic narratives, more representative of real-world use. We deliberately avoid open-ended tasks like summarisation due to unreliable evaluation metrics (Yen et al., 2024; Ye et al., 2024), focusing instead on structured-output tasks requiring factual answers, enabling precise evaluation via Exact Match Accuracy, Intersection-over-Union (IoU), and F1 score (all ranging from 0 to 1). These tasks are summarised in Table 2, with detailed descriptions in Section A.2 and examples in Appendix G.
Task Name Description Dispersion Scope Natural
@@ 3.4 Evaluation Settings
Our evaluation covers sequence lengths of 16k, 32k, 64k, and 128k tokens, using 100 samples per configuration. We evaluate all combinations of task, model size, sequence length, and sparse attention pattern at fixed compression ratios spanning 1$\times$ to 20$\times$, interpolating performance at intermediate points. We ensure input samples are 95–100% of the target maximum token length, providing a consistent basis for evaluating the impact of sequence length on performance. Following Karpinska et al. (2024), we adopt a structured prompt format (see Appendix E), encouraging models to explicitly reason via a chain of thought before providing the final answer.
@@ 4 Results
4.1 IsoFLOPS Analysis
(a) IsoFLOPS analysis for prefilling, using Vertical-Slash pattern.
(b) IsoFLOPS analysis for decoding, using Quest pattern.
Figure 2: Performance comparison for batch size 1 across FLOPS, which are a function of sequence length, model size and sparsity level. We report 4 model sizes (markers) and compression ratios up to 20$\times$ (heatmap). Performance scores are aggregated across all 9 tasks. In the plots, we display two sequence lengths—32k (left) and 128k (right)—and two phases—prefilling (top) and decoding (bottom). Crucially, there is a phase transition where after a critical sequence length (32–64k tokens for Qwen family models), highly sparse and large models surpass dense and small models in performance for the same FLOPS budget. See Appendix D for details on how we estimate the FLOPS, including indexing costs for sparse attention methods.
RQ1: Given a fixed computational budget, should one choose a small model with denser attention, or a large model with sparser attention?
Results in Figure 2 compare average performance across tasks against FLOPS for different model sizes and compression ratios at sequence lengths 32k and 128k tokens. [^4] We approach this question using the Vertical-Slash pattern for prefilling and Quest pattern for decoding, as these are, on average, the best-performing patterns for their respective inference phases (see Section 4.3). We observe a crucial effect of sequence length on the accuracy–FLOPS trade-off.
With shorter contexts (32k tokens; Figures 2(a) left and 2(b) left), most model size–sparsity combinations lie on the Pareto frontier. In this setting, decreasing sparsity yields better performance more efficiently than increasing model size. This is explained by the fact that for shorter sequences, non-attention components (embedding, MLP, and output layers) dominate the computational cost (see Appendix D for an explanation). Despite testing compression ratios up to 20$\times$, we observe that FLOPS ranges of each model size do not cross at this sequence length; the computational cost of a model twice as large (e.g., 14B vs. 7B) always exceeds the combined cost of MLP and dense attention in the smaller model, regardless of sparsity level.
@@ 4.1 IsoFLOPS Analysis
With longer contexts (128k tokens), the computational dynamics change notably, with the impact of attention sparsity becoming more pronounced. Figures 2(a) right and 2(b) right illustrate that mostly only high-sparsity models now lie on the Pareto frontier, revealing an efficiency crossover where larger sparse models surpass smaller dense ones for the same computational budget. For prefilling, models with 5–15$\times$ compression ratios remain optimal, while those with 20$\times$ compression fall below the optimal boundary. Decoding shows better resilience to high sparsity, with even 20$\times$ compression configurations being preferable to smaller models.
Looking more closely at model-specific sensitivity to sparsity, we observe distinct scaling behaviours across different model sizes. For decoding with the Quest pattern, performance resilience to sparsity correlates strongly with the model scale. Qwen 7B exhibits dramatic performance degradation under high compression, dropping by 40% (0.4 to 0.24) at 32k tokens and 79% (0.29 to 0.06) at 128k tokens when comparing dense to 20$\times$ compressed variants. In contrast, larger models (32B, 72B) maintain performance stability with compression, showing gaps consistently below 0.05 even at 20$\times$ compression. Prefilling exhibits markedly different behaviour, with performance gaps between dense and highly sparse models remaining relatively uniform (0.1–0.15) across all model sizes.
4.2 Maximum Sparsity with Guaranteed Performance
RQ2: What is the highest achievable sparsity that guarantees performance is preserved?
Figure 3: Maximum compression ratio with statistically significant performance retention (y-axis) across different model sizes (colours) and sequence lengths (x-axis). Each point represents a task, with horizontal bars showing the average maximum compression across tasks and vertical bars indicating standard deviation. Left: Vertical-Slash pattern for prefilling. Right: Quest pattern for decoding. The key conclusion is that decoding tolerates higher compression than prefilling on average, with larger models maintaining performance even at very high compression ratios. However, almost every configuration has at least one task where maximum tolerable compression is below 5$\times$ (72B Quest being the only exception).
[…]
Figure 3 reveals distinct patterns for prefilling and decoding phases. For prefilling with Vertical-Slash, we observe no clear correlation between model size, sequence length, and maximum tolerable compression. Across all sequence lengths, models can achieve compression greater than 10$\times$ without significant performance degradation when averaged across tasks. The high standard deviation (approximately 7) indicates substantial task-dependent variability in compression tolerance.
For decoding with Quest, model size emerges as a critical factor. In fact, the 7B model exhibits high variability and a clear negative correlation between sequence length and maximum compression—starting at 12$\times$ for 16k tokens but declining to 5$\times$ at 128k tokens. In contrast, larger models (32B and 72B) maintain consistently high compression ratios (around 17$\times$) regardless of sequence length. For these larger models, the majority of tasks can tolerate the maximum tested compression ratio of 20$\times$ without significant degradation.
A crucial observation is that despite these promising averages, almost every configuration has at least one task where the maximum tolerable compression is lower than 5$\times$, with 72B using Quest being the only exception (minimum 10$\times$). This highlights an important limitation: while sparse attention methods appear highly effective on average, specific inputs in each configuration may still be disproportionately affected by sparsification.
4.3 Results in Individual Tasks and Methods
@@
Figure 4: Performance comparison of different sparse attention methods across 9 tasks, aggregated over sequence lengths and models (shaded areas indicate the standard error). Top: prefilling methods (Vertical-Slash, FlexPrefill, Block-Sparse). Bottom: decoding methods (SnapKV, Ada-SnapKV, Quest). Each subplot shows the relationship between performance and compression for a specific task. The trade-off appears extremely task-dependent. Overall, Vertical-Slash performs best among prefilling methods, while Quest performs best among decoding methods.
RQ3: Are there one-size-fits-all methods consistently superior across diverse long-sequence tasks?
In Figure 4, we examine at a finer granularity how task-specific characteristics interact with various sparsification patterns across prefilling and decoding. We observe distinct trends depending on the task’s characteristics along two axes defined in our experimental setup: Scope (how much context needs to be accessed) and Dispersion (how spread out the relevant information is). We first discuss performance trends based on these task characteristics (Sections 4.3.1 and 4.3.2), then analyse the specific impact of non-uniform budget allocation strategies (Section 4.3.3), and finally summarise the findings to determine if any method is universally superior (Section 4.3.4).
4.3.1 Performance on Retrieval Tasks (Low Scope, Low Dispersion)
Retrieval tasks (QuALITY, SQuAD, TOEFL, Ruler NIAH, Story Retrieval) are characterised by Low Scope and Low Dispersion, typically requiring retrieval of specific facts from the context. For these, sparse attention methods show a clear pattern: the severity of degradation correlates directly with the number of queries. [^6] While more queries increase the required context, Retrieval tasks remain Low Scope compared to Aggregation tasks that necessitate processing the entire input. In fact, single-query QA datasets (QuALITY, SQuAD, TOEFL) maintain dense-level accuracy even at 20$\times$ compression. Ruler NIAH (4 queries) shows a slight decline for pre-filling methods and a moderate decline for decoding methods, while Story Retrieval (16 queries) experiences more substantial degradation for all methods except Quest.
Comparing units of sparsification, for methods that globally select individual tokens (Vertical-Slash, FlexPrefill, SnapKV, and Ada-SnapKV), performance drops linearly with compression ratio in the Story Retrieval task with 16 queries—a phenomenon previously studied by Yang et al. (2024c) and Chen et al. (2024). These methods estimate a key token’s importance by averaging attention scores across a subset of query tokens; however, this selection becomes harder with more queries and with fewer retrievable keys.
For block-level units of sparsification, we observe contrasting behaviour between prefilling and decoding. During prefilling, Block-Sparse performs poorly on multi-query retrieval tasks due to two key limitations (which are however necessary for efficiency reasons). First, it must process chunks of query tokens together (potentially containing multiple questions), which can lead to suboptimal importance estimation when queries are fragmented across blocks or multiple queries are collated within a single block. Second, Block-Sparse selects contiguous blocks of key tokens rather than individual tokens, introducing redundancy when retrieving factoid information that may only require specific, non-contiguous tokens.
In contrast, the chunk-level Quest pattern is the only method that maintains close-to-dense performance during decoding for Story Retrieval, even at 20$\times$ compression. Unlike Block-Sparse, Quest selects chunks of key tokens for each individual query token, offering significantly higher flexibility. Nevertheless, despite its excellent performance on Story Retrieval, Quest’s performance on Ruler NIAH drops linearly with compression, performing on par or worse with Ada-SnapKV. This degradation can be attributed to the synthetic nature of the Ruler NIAH task, which consists primarily of random symbol sequences lacking underlying syntax or semantics—leading to different distributions of key representations compared with natural language (Liu et al., 2024). This fundamental difference affects the ability of both Ada-SnapKV and Quest to differentiate between unrelated token sets, with Quest being further disadvantaged due to its coarser granularity.
4.3.2 Performance on Tasks with High Scope or High Dispersion
For tasks characterised by High Scope or High Dispersion (Ruler CWE and VT, Story Filtering and Multi-hop), which require accessing broader context or integrating information spread across the sequence, chunk-level methods (Block-Sparse, Quest) consistently match or outperform token-level global-selection methods. Specifically, Block-Sparse uniquely maintains near-baseline performance on Ruler VT at moderate compression (5–10$\times$), whereas other methods degrade significantly even at modest compression (3$\times$). Similarly, Quest notably retains high performance on Ruler CWE during decoding. We believe that these results stem from fundamental differences in how information must be processed: while attention distributions for retrieval tasks are concentrated on local windows and universally important tokens, those for tasks requiring aggregation or reasoning are more uniform (Liu et al., 2025c; Chen et al., 2024), and hence require more flexible selection mechanisms.
During prefilling, the superior performance of Block-Sparse on tasks like Ruler VT or Story Filtering can be explained by the task structures: resolving independent chains of variables or examining multiple chapters independently is better suited for chunk-based methods compared with global token selection methods, which struggle to assign their limited budget in these tasks. For decoding, Quest outperforms on complex tasks by preserving the complete key-value cache while only optimising memory transfers, unlike SnapKV variants which irreversibly prune tokens before generation begins. Crucially, this severely limits the model’s ability to perform inference-time scaling through strategies like chain-of-thought reasoning as these rely on the ability to explore alternative reasoning paths—and hence possibly different sets of important tokens (Li et al., 2025b, 2024b)—when initial approaches prove incorrect (DeepSeek-AI, 2025).
4.3.3 Impact of Non-Uniform Budget Allocation
Non-uniform budget allocation strategies yield distinct outcomes depending on the inference phase. During prefilling, FlexPrefill’s adaptive, threshold-based selection either matches or underperforms Vertical-Slash’s simpler, uniform allocation across tasks. Ablations (Section A.1) revealed FlexPrefill’s extreme sensitivity to the minimum-budget hyperparameter: low values can degrade performance, while high values effectively mimic Vertical-Slash’s fixed-budget scheme at higher compression ratios. This behaviour echoes the “attention sink” phenomenon (Chen et al., 2024), where threshold-based methods might capture a few high-attention tokens, thus being satisfied with their cumulative attention recall criteria, while missing critical information distributed across the long tail of the attention distribution. Conversely, during decoding, Ada-SnapKV consistently outperforms the uniform SnapKV, particularly for complex multi-query retrieval tasks that inherently require broader token coverage, demonstrating the benefit of adaptive allocation when generation depends on accessing diverse context elements.
4.3.4 Summary and Method Recommendations
Our comprehensive evaluation indicates no single sparse attention method uniformly excels across all diverse long-sequence tasks, although methods offering greater flexibility generally perform better. For prefilling, Vertical-Slash consistently delivers optimal results on retrieval tasks (Low Scope, Low Dispersion), yet slightly trails behind chunk-level methods like Block-Sparse on aggregation and multi-hop reasoning tasks (High Scope or Dispersion). For decoding, Quest emerges as the most robust approach due to its high flexibility in selecting chunks per query and maintaning access to the entire KV cache, particularly adept at handling complex tasks; however, it remains vulnerable to synthetic tasks like Ruler NIAH and incurs additional memory overhead for maintaining page-level summaries required for its efficient selection mechanism. Therefore, the optimal choice of sparse attention method is highly dependent on the specific task characteristics, the required flexibility, and the inference phase (prefilling vs. decoding).
4.4 Sparse Attention Scaling Laws
RQ4: Can we establish scaling laws for sparse attention, which can generalise to larger model sizes, sequence lengths, or compression ratios?
We aim to fit scaling laws that predict downstream accuracy given task, model size, sequence length, and sparse attention compression ratio. We focus on the best-performing sparse attention pattern for each generation phase, namely Vertical-Slash for sparse pre-filling and Quest for sparse decoding.
After surveying previous scaling laws studies in foundation models, an overview of which is available in Appendix B, we opt for a log-linear scaling law as it makes the simple assumption of a power law with task-specific intercepts. Thus, our scaling law takes the form of
$s=\alpha\log N+\beta\log L+\gamma\log C+\delta_{task}+\epsilon,$
where $s$ is the logit (inverse sigmoid) of the accuracy, $N$ is model parameter count, $L$ is sequence length, $C$ is compression ratio, $\delta_{task}$ is a task-specific intersect and $\epsilon$ is a shared intercept. We provide more details on our scaling law formulation and fitting process in Appendix C.
We perform the fitting and validation of our scaling law on the Qwen 2.5 model family. We conduct 3 runs, each one holding out data points with the highest value for one axis as the test set, in order to evaluate the scaling law’s ability to extrapolate along that specific axis. Concretely, we hold out data points with the largest model ($72$B), the longest sequence ($128$k) and the highest compression ratio (20×20$\times$20 ×) respectively in each run.
In terms of fitting, we choose $R^{2}$ as the main metric, as it measures how much of the total variance is explained by our scaling law model. We report $R^{2}$ and the inferred parameters in Table 3. From Table 3 it emerges that $R^{2}$ is very high (between 0.57 and 0.74). This demonstrates that the fit is very robust and our choice of predictors is largely appropriate for capturing the underlying pattern in the data. What is more, the parameters reveal that task-specific intersects mirror our task categorisation: their value is negative for most multi-hop and aggregation tasks (Ruler VT and Story Filtering and Multi-Hop). The intercept for Ruler NIAH stands out for its exceptionally high value.
$\mathbf{R^{2}}$ Parameters
$\epsilon$ $\delta_{\text{QA}}$ $\delta_{\text{Ruler}}$ $\delta_{\text{Story}}$ $\alpha$ $\beta$ $\gamma$
SQuAD TOEFL CWE NIAH VT Filt. Mult. Retr.
Prefill Sequence Length 0.74 -16.92 0.27 0.80 0.76 14.67 -1.60 -3.55 -3.21 1.43 1.31 -1.15 -1.00
Model Size 0.64 -18.10 0.20 0.73 -0.42 10.51 -1.63 -4.42 -3.74 0.46 1.58 -1.58 -0.83
Compression Ratio 0.72 -18.08 0.24 0.79 0.40 13.62 -1.54 -3.65 -3.55 1.17 1.51 -1.51 -0.73
Decode Sequence Length 0.64 -28.82 0.35 0.72 0.54 12.18 -0.85 -2.98 -2.87 2.32 1.72 -0.97 -0.94
Model Size 0.57 -35.57 0.32 0.60 -0.53 8.45 -0.89 -3.88 -3.44 1.31 2.25 -1.45 -0.81
Compression Ratio 0.64 -26.89 0.31 0.66 0.26 11.37 -0.90 -3.19 -3.40 1.78 1.81 -1.38 -0.82
Table 3: $R^{2}$ as a measure of fitness and inferred parameters for our scaling laws.
To evaluate the quality of the scaling laws’ predictions, for each of the three axes of extrapolation and for each of the nine tasks, we report two main metrics on test examples: 1) the Spearman’s $\rho$ between true and predicted values; and 2) the mean absolute error (MAE) of the predicted accuracy. The evaluation results of our scaling laws are shown in Table 4 in terms of Spearman’s $\rho$ and in Table 7 in Section C.3 in terms of MAE.
QA Ruler Story
Quality SQuAD TOEFL CWE NIAH VT Filtering Multi-hop Retrieval
Prefill Sequence Length .73 .79 .78 .86 .87 .86 .90 .34 .82
Model Size .68 .75 .60 .85 .56 .68 .89 .90 .85
Compression Ratio .77 .87 .63 .75 .82 .40 .79 .38 .65
Decode Sequence Length .75 .81 .69 .93 .84 .54 .67 .50 .92
Model Size .77 .49 .31 .92 .52 .84 .84 .89 .81
Compression Ratio .88 .91 .87 .97 .75 .92 .96 .59 .93
Table 4: Spearman’s correlation $\rho$ by task between true and predicted accuracy on each held-out axis (sequence length, model size, compression ratio) and each phase (prefilling and decoding). We mark statistically significant entries where $p<0.05$ with .
Based on Table 4, we find that our scaling laws can meaningfully predict downstream accuracy in new configurations, as correlations between true and predicted performance are generally strong. Nonetheless, there remain a few settings where the correlation is weak or not statistically significant. Most notably, this is the case for the task of Story Multi-hop, where our laws fail to extrapolate accuracy when ablating the compression ratio (during both phases) and sequence length (during prefilling).

### v2

@@ 3.1 Models
We perform experiments primarily on Qwen 2.5 (Yang et al., 2024a) (7B, 14B, 32B, 72B parameters), complemented by Llama 3.1 (Dubey et al., 2024) (8B, 70B) and Gemma 3 (Gemma Team, 2025) (4B, 12B, 27B). All three families use instruction-tuned variants to support chain-of-thought evaluation. Qwen 2.5 was selected as our primary family as it uniquely satisfies strict methodological requirements for controlled scaling experiments—see Section A.3 for rationale. For Qwen and Llama, we modify the attention mechanism across all layers. Gemma 3 employs hybrid attention where 5 out of 6 layers use sliding window attention (1024 tokens) by design; we apply sparse attention methods only to the remaining dense (global attention) layers. We preserve the original architectures and utilise the vLLM inference engine (Kwon et al., 2023) with full bf16 precision. Implementation details are in Section A.1.
@@ 3.2 Sparse Attention Methods
We evaluate six state-of-the-art sparse attention methods (Table 1), which we choose as a representative set spanning across the key dimensions described in Section 2. We focus exclusively on content-aware methods, as prior work has demonstrated that fixed patterns consistently underperform their dynamic counterparts (Li et al., 2025).
@@ 3.3 Tasks
We evaluate 9 diverse tasks selected to reflect different characteristics along 3 key dimensions known to influence sparse attention performance: task difficulty—defined by Dispersion (how hard it is to locate necessary information) and Scope (how much information must be processed) (Goldman et al., 2024)—and data Naturalness (natural language vs. synthetic data). This multi-dimensional approach is motivated by recent findings that attention patterns vary significantly across task types: retrieval tasks often exhibit localised attention, while reasoning tasks show more uniform distributions that are challenging for sparse methods (Liu et al., 2025c; Chen et al., 2024; Li et al., 2025). The naturalness dimension is also crucial, as synthetic tasks yield different token representation distributions compared to natural language (Liu et al., 2024). Our task suite therefore incorporates four core tasks from the RULER benchmark (Hsieh et al., 2024)—Retrieval (NIAH), Multi-hop reasoning (VT), Aggregation (CWE), and QA (SQuAD)—to provide controlled environments (mostly synthetic) for specific capabilities. We complement these with natural texts from benchmarks with minimal contamination risk (Li et al., 2024a), such as QuALITY and TOEFL, though these represent low-dispersion, low-scope tasks. Thus, we additionally introduce three novel tasks (Story Retrieval, Multi-hop, Filtering) that translate RULER’s challenging tasks (with high dispersion or scope) into naturalistic narratives, more representative of real-world use. We deliberately avoid open-ended tasks like summarisation due to unreliable evaluation metrics (Yen et al., 2024; Ye et al., 2024), focusing instead on structured-output tasks requiring factual answers, enabling precise evaluation via Exact Match Accuracy, Intersection-over-Union (IoU), and F1 score (all ranging from 0 to 1). These tasks are summarised in Table 4, with detailed descriptions in Section A.2 and examples in Appendix G.
@@ 3.4 Evaluation Settings
Our evaluation covers input lengths of 16k, 32k, and 64k tokens for all model families, with 128k evaluations limited to Qwen and Llama using Vertical-Slash and Quest only; Gemma exhibited near-zero performance at 128k. We use 100 samples per configuration for Qwen and 50 for Llama and Gemma. We evaluate all combinations of task, model size, sequence length, and sparse attention pattern at sparsity levels from 0 (dense) to 0.95 (i.e., $1/20$ attention budget), interpolating performance at intermediate points. We ensure input samples are within 95–100% of the target maximum token length, providing a consistent basis for evaluating the impact of sequence length on performance. Following Karpinska et al. (2024), we adopt a structured prompt format that encourages models to explicitly reason through chain-of-thought before providing answers in a consistent, parsable structure (see Appendix E). As metrics of computational cost, we report FLOPS for prefilling and memory access for decoding, as these reflect the respective computational bottlenecks of each phase (see Section 2). Appendix B provides more details, including indexing costs for sparse attention methods.
@@ 4 Results
4.1 isoCost Analysis
Figure 1: isoCost analysis for Qwen 2.5 (128k tokens) and Gemma 3 (64k tokens). Each point corresponds to a (model size, sparsity) configuration, with performance aggregated across 9 tasks. Left column: prefilling with Vertical-Slash (Jiang et al., 2024) (FLOPs). Right column: decoding with Quest (Tang et al., 2024) (memory transfers). Standard error is negligible (Section D.1) and omitted for visual clarity. Dotted lines show Pareto frontiers connecting configurations that are not dominated by any other configuration. Key findings: (1) sparsification enables larger sparse models to outperform smaller dense models at equivalent cost; (2) the impact of sparsity is less pronounced for Gemma due to its sliding-window architecture (Figure 14).
RQ1: Is sparse attention effective? Results in Figure 1 illustrate the average performance across tasks against computational cost for different model sizes and levels of sparsity.We approach this question using Vertical-Slash for prefilling and Quest for decoding, as these are, on average, the best-performing patterns for their respective inference phases (see Section 4.2). As implementation-agnostic proxies for computational cost, we use FLOPs for prefilling and memory transfers for decoding, which correlate with wall-clock time under optimised implementations (see Appendix B for cost formulas and breakdowns). We visualise Pareto frontiers to identify configurations that offer the best performance-cost trade-offs, i.e., those not dominated by any other configuration in terms of both cost and performance.
Sparse attention improves the Pareto frontier.
@@
In Figure 1, the Pareto frontier reveals an efficiency crossover where sparsification enables larger sparse models to outperform smaller dense ones at equivalent computational cost. For Qwen at 128k tokens, only high-sparsity configurations lie on the Pareto frontier. During prefilling, models with sparsity 0.8–0.93 (i.e., $1/5$ to $1/15$ attention budget) remain optimal, while sparsity 0.95 ($1/20$ budget) falls below the optimal boundary. Decoding shows better resilience to high sparsity, with even 0.95 sparsity configurations being preferable to smaller dense models. For Gemma, we observe similar trends during decoding, but configuration overlap is absent for prefilling—this reflects Gemma’s lower baseline attention ratio because of its sliding-window architecture (see Sections 2 and B).
4.2 Per-Task Analysis
RQ2: Which sparse attention method should practitioners use? Figure 2 presents per-task performance across sparse attention methods, aggregated over three model families and sequence lengths up to 64k. The 9 tasks introduced in Section 3.3 are grouped by their information retrieval characteristics: Single QA (one query, localised answer), Multiple QA (multiple queries targeting distinct facts), High Scope/Low Dispersion (broad context, concentrated answers), and Low Scope/High Dispersion (narrow focus, scattered information). Three findings emerge from this analysis.
@@ 4.2 Per-Task Analysis
Figure 2: Per-task performance comparison of sparse attention methods, aggregated over Qwen 2.5, Llama 3.1, and Gemma 3 models at sequence lengths 16k, 32k, and 64k. Error bars indicate standard error. Left column: prefilling (Vertical-Slash, FlexPrefill, Block-Sparse). Right column: decoding (SnapKV, Ada-SnapKV, Quest). Tasks are grouped by information retrieval characteristics. Per-family breakdowns are provided in Section D.2.
Prefill and decoding phases display different flexibility.
As discussed in Section 2, the computational constraints of each inference phase fundamentally determine what patterns can be selected, which in turn affects generalisation across tasks. During prefilling, per-cell importance estimation would require quadratic cost, leaving only two strategies: global selection of fine-grained units (Vertical-Slash) or block-to-block selection (Block-Sparse). Neither strategy dominates—the optimal choice is task-dependent.
During prefilling, Vertical-Slash shows strong performance on retrieval tasks (Low Scope, Low Dispersion) by enabling fine-grained token selection for locating specific facts. Tasks demanding broader context access or multi-step reasoning (High Scope or Dispersion, e.g., Ruler VT, Story Filtering) benefit from Block-Sparse, which selects distinct key-token blocks for each query block, accommodating the processing of multiple independent segments.
During decoding, token-to-page selection becomes computationally feasible since only one query is processed per step. This greater flexibility enables Quest to generalise better across tasks and tolerate higher compression than either prefilling approach, while retaining the full KV cache. Eviction-based decoding methods (SnapKV, Ada-SnapKV) that permanently discard tokens illustrate the cost of sacrificing the full cache—irreversible compression is detrimental when discarded tokens become relevant later, though this comes with the benefit of reduced memory footprint. Nevertheless, Quest can degrade on synthetic tasks such as Ruler NIAH, where random symbol sequences yield less distinguishable key representations compared to natural language (Liu et al., 2024)—Quest’s page-level granularity amplifies this effect, as coarser blocks struggle more than Ada-SnapKV’s token-level selection to differentiate between unrelated token sets.
Dynamic budget allocation benefits are phase-dependent.
Adaptive methods that allocate different budgets across layers or sequences yield inconsistent results. During prefilling, FlexPrefill matches or underperforms Vertical-Slash’s uniform allocation, likely due to the “attention sink phenomenon” (Chen et al., 2024): threshold-based selection captures high-attention tokens while missing information in the distribution’s long tail. During decoding, Ada-SnapKV consistently outperforms uniform SnapKV, particularly on multi-query tasks (Story Retrieval), though both eviction methods remain inferior to Quest’s full-cache approach.
Sparsity tolerance varies dramatically across tasks.
The gap between task groups reveals a deployment risk: methods achieving high sparsity on easy tasks may fail on harder ones. Single QA tasks (QuALITY, SQuAD, TOEFL) tolerate sparsity 0.95 ($1/20$ budget) with minimal degradation across all methods. Multiple QA tasks (Ruler NIAH, Story Retrieval) show substantial degradation at sparsity 0.8–0.9 ($1/5$ to $1/10$ budget). Tasks with High Scope or High Dispersion degrade even at modest sparsity (0.5–0.67, i.e., $1/2$ to $1/3$ budget) for some methods. Evaluating sparse attention only on Single QA benchmarks—or averaging across task types—masks these vulnerabilities. Robust deployment requires testing across diverse task characteristics, as sparsity levels safe for retrieval tasks can cause failures on aggregation or multi-hop reasoning. Moreover, sequence naturalness affects methods asymmetrically—Quest outperforms Ada-SnapKV on natural-language retrieval (Story Retrieval) but underperforms on synthetic retrieval (Ruler NIAH)—underscoring the need for benchmarks spanning both natural and synthetic data.
4.3 Sequence Length Effects
Figure 3: Sequence length effects on sparsity tolerance. Relative error is $(\bar{p}_{\text{dense}}-\bar{p}_{\text{sparse}})/\bar{p}_{\text{dense}}$, where $\bar{p}$ denotes mean performance. Results aggregated across all tasks, methods, and models (Qwen 2.5, Llama 3.1, Gemma 3). Per-family breakdowns are provided in Section D.3.
RQ3: How does sequence length affect tolerance to sparse attention? Figure 3 shows that for a fixed attention budget fraction, longer sequences incur smaller degradation: for example, at a $1/20$ budget, the relative error decreases from $\approx 0.33$ (16k) to $\approx 0.26$ (32k) and $\approx 0.20$ (64k). This indicates that the same sparsity ratio becomes less harmful as the sequence length grows. This pattern holds consistently across all model families. Nawrot et al. (2024) observe similar results for a training-aware KV compression method: their learned mechanism applies lower sparsity at the beginning of sequences and increases sparsity with sequence length. This behaviour may be explained by Herdan’s law (Herdan, 1960), which posits that new information becomes rarer over time, facilitating higher sparsity with distance.
To relate this trend to budget scaling, we interpret the plot as approximate iso-error curves. For a target relative error of $\approx 0.2$, the required budget fractions are roughly $1/10$ (16k), $1/15$ (32k), and $1/20$ (64k). In contrast, a fixed token budget would imply fractions $1/10\rightarrow 1/20\rightarrow 1/40$ as length grows, which already exceeds the $\approx 0.2$ error target at 32k (the $1/20$ point is $\approx 0.26$). For a stricter target such as $\approx 0.1$, the scaling is less uniform: $1/5$ stays below $0.1$ for both 16k and 32k, while 64k requires only a modest reduction in sparsity to stay near $0.1$. These observations imply that the optimal token budget should grow sublinearly with sequence length: doubling the context does not require doubling the token budget, but keeping the budget constant would incur increasing degradation. While current dynamic methods lack robustness (Section 4.2), developing reliable sublinear budget allocation mechanisms remains a promising direction for future work.